26 is what % of 20
Answer by ptaylor(2198)

First we'll let x equal the fractional part of 20 that 26 happens to be and then we will multiply that by 100% to get the percent:
soooo:
20*x=26
x=26/20=1.3
1.3*100%=130%-------------answer
CK
20*1.3=26
26=26
Hope this helps---ptaylor